OpenStudy (anonymous):

Solve the system by the elimination method. x + y - 6 = 0 x - y - 8 = 0 When you eliminate y , what is the resulting equation?

OpenStudy (tkhunny):

Have you considered adding the two equations - combining like terms?

OpenStudy (anonymous):

x+-6=0 x+8=0

OpenStudy (tkhunny):

@alex1248191 Ummm.... No. Please reconsider your response.

OpenStudy (bradely):

Add two equations 2x-14=0 x=7 x+y-6=0 7+y-6=0 y=-1 Source: http://www.mathskey.com/question2answer/

OpenStudy (anonymous):

dang, my bad

OpenStudy (anonymous):

(1) x + y - 6 = 0 (2) x - y - 8 = 0 _____________________________________ (1)+(2)--->2x-14=0 x=7; (1)-(2)---->2y+2=0 y=-1.